[figure: 9.7 cm]

QUANTUM CRITICAL BEHAVIOR IN CuGeO_{3}

S.V.Demishev^{a}, Y.Inagaki^{b}, M.M.Markina^{c}, H.Ohta^{b}, S.Okubo^{b}, Y.Oshima^{b}, A.A.Pronin^{a}, N.E.Sluchanko^{a}, N.A.Samarin^{a} aGeneral Physics Institute of RAS, Vavilov street, 38, 119991 Moscow, Russia bKobe University, 1-1 Rokkodai, Nada, Kobe 657-8501, Japan cMoscow State University, 119899 Moscow, Russia

High frequency ESR in the spin-Peierls compound CuGeO_{3} doped with 1% Fe and 2% Co was studied. Experiments were carried out in the frequency/temperature/field domain 60-450 GHz/1.8-300 K/0-15 T. The use of the ESR spectroscopy have allowed to separate correctly effects of magnetism of Cu^{2+} chains modified by magnetic impurity from Fe or Co impurity paramagnetism and found magnetic susceptibility of Cu^{2+} chains χ(T), line widths and g−factors. It was found that doping of CuGeO_{3} matrix by Fe and Co completely damps both spin-Peierls and antiferromagnetic transitions. For T <20K a divergent power asymptotic χ∝1/T^{α} with the indexes α=0.35±0.03 (CuGeO_{3}:Fe) and α=0.69±0.04 (CuGeO_{3}:Co) was observed. Simultaneously for CuGeO_{3}:Fe magnetic contribution to specific heat for T <20 K acquires the form C_{m} ∝ T^{1−α} with α=0.37±0.04. The observed anomalies suggest onset of the quantum critical behavior^{1} in CuGeO_{3} doped with magnetic impurities.

1 Demishev S.V., et al. JETP Letters 73, 31 (2001)

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : oral

Corresponding author : S.V.Demishev

Address for correspondence : General Physics Institute of RAS, Vavilov street, 38, 119991 Moscow, Russia

Email address : demis@lt.gpi.ru